OpenStudy (anonymous):

Help Please!!! Solve agebraically. Answer in interval notation (x^3)-(4x^2)>0

OpenStudy (anonymous):

start with \[x^2(x-4)>0\] and then it should be real easy, since \(x^2\geq 0\) for all \(x\)

OpenStudy (anonymous):

oh that's great! so what is the answer you think? I think i got (infinity,4)....;;; not sure

OpenStudy (anonymous):

well, there is no such thing as \((\infty, 4)\) that is for sure

OpenStudy (anonymous):

try \[(4,\infty)\]

OpenStudy (anonymous):

or it could be (-infinity,4)?

OpenStudy (anonymous):

well im stucked again;;;

OpenStudy (anonymous):

no it cannot be \((-\infty, 4)\) since you are solving \[x-4>0\]

OpenStudy (anonymous):

so the final answer could be (4,infinity)?